# Model Alignment Search

### v6

@@ Abstract
When can we say that two neural systems are the same? The answer to this question is goal-dependent, and it is often addressed through correlative methods such as Representational Similarity Analysis (RSA) and Centered Kernel Alignment (CKA). What nuances do we miss, however, when we fail to causally probe the representations of the systems? Do the dangers of cause vs. correlation exist in comparative representational analyses? In this work, we introduce a method for connecting neural representational similarity to behavior through causal interventions. The method learns orthogonal transformations that find an aligned subspace in which behavioral information can be isolated and interchanged between multiple distributed networks’ representations. We first show that the method can be used to transfer the behavior from one frozen Neural Network (NN) to another in a manner similar to model stitching, and we show how the method can differ from correlative similarity measures like RSA. Next, we empirically and theoretically show how our method can be equivalent to model stitching when desired, or it can take a form that is more restrictive, narrowing its focus to shared causal information; and in both forms, it reduces the number of required matrices for a comparison of n models to be linear in n. We then present a case study in number-related tasks showing that the method can be used to examine specific subtypes of causal information, and we present another case study showing that the method can reveal toxicity in fine-tuned DeepSeek-r1-Qwen-1.5B models. Lastly, we show how we can augment the loss with a counterfactual latent auxiliary objective to improve causal relevance when one of the two networks during training is causally inaccessible (as is often the case in comparisons with biological networks). We use our results to encourage the use of causal methods in neural similarity analyses and to suggest future explorations of network similarity methodology for understanding LLM misalignment.
@@ 1 Introduction
An important question for understanding both Artificial and Biological Neural Networks (ANNs and BNNs) is knowing what it means for one distributed system to model or represent another (Sucholutsky et al., 2023). Establishing isomorphisms between different distributed systems can be useful for simplifying their complexity and for understanding otherwise opaque inner mechanisms. Instead of asking how to interpret neural activity directly, we can instead use simplified neural systems, or neural systems understood through a higher level of analysis, and compare the internals of these known systems to uknown systems (Cao & Yamins, 2021; 2024; Richards et al., 2019). Furthermore, there are a number of open questions about how representations differ or converge between architectures, tasks, and modalities (Huh et al., 2024; Sucholutsky et al., 2023; Wang et al., 2024; Li et al., 2024; Hosseini et al., 2024; Zhang et al., 2024; Grant et al., 2025). Implicitly intertwined in these issues is the notion of network similarity, or dissimilarity. How do we measure the degree to which one model represents another?
@@
Figure 1: (a) A depiction of an interchange intervention from Equation 6 on the target latent vector $h^{(trg)}_{\psi_{i}}$ from model $\psi_{i}$ using the source latent vector $h^{(src)}_{\psi_{k}}$ from $\psi_{k}$. Rectangles represent vectors; colors distinguish between behaviorally relevant and extraneous activity. The causally relevant information is spread across the neural population (Smolensky, 1988; Park et al., 2023), represented by the red and green semi-vertical slices in the respective $h$ vectors. To disentangle and isolate the complement of the behavioral null-space, we rotate the $h$ vectors into an aligned space, using $Q_{\psi_{i}}$ and $Q_{\psi_{k}}$, where the behaviorally relevant information is organized along separate dimensions than the extraneous, behavioral null-space. We can then intervene on and transfer the relevant information without affecting other information. In the figure, this is done by applying binary masks (black represents 0s and white 1s and $\odot$ is a Hadamard product) to the $z$ vectors and taking their sum. We then invert the rotation on $z^{(v)}_{\psi_{i}}$ to return it to the original latent space where it can be used by $\psi_{i}$ to make predictions. (b) Depicts Stepwise MAS where the individual intervention shown in panel (a) is applied at multiple token positions in the sequence. We limit our interventions to contiguous sets of tokens starting from the first token and ending with a sampled position $t$.
@@
Researchers often use correlational methods to measure the similarity of different neural representations. We can see examples of this in works that perform direct correlational analyses between individual ANN activations and BNN firing rates (Yamins & DiCarlo, 2016; Maheswaranathan et al., 2019; Khosla & Williams, 2023; Williams et al., 2022), and in works that use Representational Similarity Analysis (RSA)—or Centered Kernel Alignment (CKA) (Kornblith et al., 2019; Williams, 2024)—finding 2nd order isomorphisms between model and system (Kriegeskorte et al., 2008). We also see examples of this in linear decoding techniques, where linear decodability and predictability can be used as a measure of the type of information encoded in distributed representations (Chen et al., 2020; Radford et al., 2021; Grill et al., 2020; Caron et al., 2021; Haxby et al., 2001; Haxby, 2013; Lampinen et al., 2024). A question remains, however, how to causally associate these neural analyses with behavioral outcomes. What do we miss when we ignore the behavioral relevance of the neural activity (Pearl, 2010; Geiger et al., 2024; Cloos et al., 2024; Schaeffer et al., 2024)? And how do we incorporate a focus on behavioral relevance in our measures of neural similarity?
Some works have made progress towards connecting representational similarity to behavior by transforming intermediate representations from one neural system into a usable form for another. We see examples in works like Sexton & Love (2022) where they attempt to use transformed neural recordings in a trained computational model, and in model stitching, where a linear mapping is learned from intermediate representations from one ANN to another for the purpose of measuring similarity and/or improving one of the two models (Lähner & Moeller, 2024; Moschella et al., 2023; Bansal et al., 2021; Lenc & Vedaldi, 2015). To build upon these works, we ask: 1) what do these representational mappings tell us about the underlying representations of the two systems; are behaviorally successful mappings both necessary and sufficient for claims of functional similarity? 2) How can we isolate and compare the behaviorally relevant activity within transferred representations? 3) How do we isolate the similarity of specific behavioral information in the representations, thus allowing us to compare models with disparate behavioral outputs? And 4) how do we achieve causal relevance in systems for which we do not have causal access—as is often the case in ANN to BNN comparisons?
@@
In this work, we introduce a neural similarity method, Model Alignment Search (MAS), that uses causal interventions to isolate and compare behaviorally relevant activity from neural representations in different ANNs. MAS learns a rotation matrix, or more generally an alignment function, for each model in the analysis to simultaneously uncover causally relevant latent subspaces in each model and map these spaces onto one another so that behavioral information can be patched between models. These manipulated representations are then returned back to their original neural space where the networks’ resulting behavior can be compared to the expected counterfactual behavior as a means of measuring neuro-behavioral similarity.
@@
    We introduce and motivate MAS by comparing to correlative representational similarity methods demonstrating a need to supplement correlational with causal methods depending on the research goals.
@@
    We introduce a counterfactual latent auxiliary loss objective that can be used to find behaviorally relevant alignments in cases where one of the two models is causally inaccessible (making the technique potentially relevant for comparisons between ANNs and BNNs).
@@ 2.1 Model Stitching
where $x~\mathcal{D}$ are samples from a dataset and $\ell$ is a divergence measure (e.g., KL divergence or cross-entropy) between $\psi_{1}$’s output distribution and $\psi_{2}$’s output distribution under the stitched representation.
@@
Model stitching has been used as a method to make claims about the similarity between two different NNs Bansal et al. (2021). However, it has problems as a similarity method due to the fact that it is unidirectional, and, as we will show in Section 4.2, it can make use of $\psi_{1}^{\prime}s$ behavioral null space when mapping to $\psi_{2}$—where the behavioral null space is defined as the span of vector directions that have no influence on the NN’s behavior. Furthermore, a successful mapping ignores naturally occurring variability in $\psi_{2}$’s behavioral null space. Thus, a successful mapping $T$ only needs to produce a single sufficient representation for any given behavior. Thus, successful $T$’s do not necessarily indicate that $\psi_{1}$ and $\psi_{2}$ perform the task in the same way.
@@ 2.2 DAS Formulation
DAS is a framework for causally determining the degree of alignment between a single NN’s latent representations and variables from a causal abstraction (Geiger et al., 2021; 2023; Wu et al., 2024). We can think of such variables as those in a computer program, or nodes in a Directed Acyclic Graph (DAG). DAS does this alignment by attempting to transform model latent states $h\in\mathbb{R}^{d_{\psi}}$ into aligned vectors $z\in\mathbb{R}^{d_{\psi}}$ that consist of contiguous subspaces that have an analogous causal effect on behavior as high-level variables from a pre-specified causal abstraction. The transformation is performed by learnable, invertible Alignment Function (AF), $\mathcal{A}$, as follows: $z=\mathcal{A}(h)$ (Grant et al., 2025). We narrow our focus to orthogonal AFs: $\mathcal{A}(h)=Qh$ where $Q\in\mathbb{R}^{d_{\psi}\times d_{\psi}}$ is an orthogonal matrix. The purpose of this transformation is that it organizes causal subspaces into separated, contiguous dimensions in $z$. This allows for causal interventions on the values of specific subspaces—analogously to changing value of a variable in a computer program—without affecting the values of the other subspaces.
@@ 3.3 MAS Formulation
With these intutions, we can generalize Equation 3 to the multi-model case for an intervention on the subspace for $\text{var}_{k}$:
$h_{\psi_{i}}^{(v)}=Q_{\psi_{i}}^{-1}((1-D_{\psi_{i},\text{var}_{k}})Q_{\psi_{i}}h^{(trg)}_{\psi_{i}}+D_{\psi_{j},\text{var}_{k}}Q_{\psi_{j}}h^{(src)}_{\psi_{j}})$ (6)
Where $i$ and $j$ can be any model index in the set of all models considered, $Q_{\psi_{i}}$ is a scaled orthogonal rotation matrix for $\psi_{i}$, $D_{\psi_{i},\text{var}_{k}}\in R^{d_{\psi_{i}}\times d_{\psi_{i}}}$ is a diagonal, binary matrix with $d_{\text{var}_{k}}$ non-zero elements used to isolate the dimensions corresponding to $\vec{z}_{\text{var}_{k}}$, $h^{(src)}_{\psi_{j}}$ is the source vector from which the subspace is harvested, $h^{(trg)}_{\psi_{i}}$ is the target vector into which activity is substituted, and $h^{(v)}_{\psi_{i}}$ is the resulting intervened vector that replaces $h^{(trg)}_{\psi_{i}}$ in $\psi^{(trg)}_{i}$’s processing, allowing the model to make causally intervened predictions. See Figure 1(a) for a visualization.
For many of our analyses, we will perform MAS without segregating the behaviorally relevant subspace. Concretely, this can be formalized as a causal abstraction that uses a single Full variable in which all behaviorally relevant information is encoded, $\vec{z}_{full}$, and all extraneous information is encoded in $\vec{z}_{extra}$, the behavioral null-space and dormant subspaces. For completion: $z_{\psi_{i}}=\begin{bmatrix}\vec{z}_{\psi_{i},full}\\ \vec{z}_{\psi_{i},extra}\end{bmatrix}$, and we will freely isolate and manipulate $\vec{z}_{\psi_{i},full}$ using Equation 3.
@@ 3.4 MAS Training
It is crucial to include interventions where $i=j$ as this adds a soft constraint to the alignment that encourages the separation of $\vec{z}_{full}$ from $\vec{z}_{extra}$.
@@ 3.5 Mas Variants
3.6 Representational Similarity Analysis (RSA)
Figure 2: (a) and (b) A comparison of MAS on the left axes and CKA and RSA on the right axes. We examine both the input embeddings and the hidden state vectors for models trained on the Multi-Object task. (a) Results for GRUs compared to GRUs, where RSA can give low estimates of embedding similarity for different model seeds whereas MAS shows the high causal transfer we might expect. (b) Results for GRUs compared to 2-layer Transformers where we see a similar effect as (a) in the embeddings using RSA and we see a potential over-estimation of similarity of the hidden states in CKA and RSA. This over-estimation is with respect to causal transfer, as prior work has shown the transformers to use anti-Markovian solutions, where they recompute the relevant information at each step in the sequence. This is reflected in the low MAS IIA (Grant et al., 2025). (c) IIA comparing finetuned toxic and nontoxic LLMs using stepwise MAS. We can see that toxic models have higher IIA with themselves than with the nontoxic models. Notably, there is no significant difference for the nontoxic models compared to themselves. (d) Comparison of the IIA from DAS and MAS for different sizes of the aligned subspace, and model stitching with different rank transformation matrices.

### v7

@@ Abstract
When can we say that two neural systems perform a task in the same way? What nuances do we miss when we fail to causally probe the representations of the systems, and how do we establish bidirectional causal relationships? In this work, we introduce a method that bidirectionally transfers neural activity between artificial neural networks and uses their resulting behavior as a measure of functional similarity. We first show that the method can be used to transfer the behavior from one frozen Neural Network (NN) to another in a manner similar to model stitching, and we show how the method can differ from correlative similarity measures like Representational Similarity Analysis. Next, we empirically and theoretically show how the method can be equivalent to model stitching when desired, or it can take a form that has a more restrictive focus to shared causal information; in both forms, it reduces the number of required matrices for a comparison of n models to be linear in n. We then present a case study on number-related tasks showing that the method can be used to examine specific subtypes of causal information demonstrating that numbers can be encoded differently in recurrent models depending on the task, and we present another case study showing that MAS can reveal misalignment in fine-tuned DeepSeek-r1-Qwen-1.5B models. Lastly, we augment the loss function with a counterfactual latent (CL) auxiliary objective to improve causal relevance when one of the two networks is causally inaccessible (as is often the case in comparisons with biological networks). We use our results to encourage the use of causal methods in neural similarity analyses and to suggest future explorations of network similarity methodology for model misalignment.
@@ 1 Introduction
An important question for understanding both Artificial and Biological Neural Networks (ANNs and BNNs) is knowing what it means for one distributed system to model or represent another (Sucholutsky et al., 2023). Establishing isomorphisms between different distributed systems can be useful for simplifying their complexity and for understanding otherwise opaque inner mechanisms. Instead of asking how to interpret neural activity directly, we can instead compare the internals of the simplified neural systems, or complex neural systems understood through a higher level of analysis (i.e. architecture, learning rule, training data), to the internals of the unknown systems (Cao & Yamins, 2021; 2024; Richards et al., 2019). In addition, there are several open questions about how representations differ or converge between architectures, tasks, and modalities (Huh et al., 2024; Sucholutsky et al., 2023; Wang et al., 2024; Li et al., 2024; Hosseini et al., 2024; Zhang et al., 2024; Grant et al., 2025).
@@
Figure 1: (a) A depiction of an interchange intervention from Equation 6 on the target latent vector $h^{(trg)}_{\psi_{i}}$ from model $\psi_{i}$ using the source latent vector $h^{(src)}_{\psi_{k}}$ from $\psi_{k}$. Rectangles represent vectors; colors distinguish between behaviorally relevant and extraneous activity. The causally relevant information is spread across the neural population (Smolensky, 1988; Park et al., 2023), represented by the red and green semi-vertical slices in the respective $h$ vectors. To disentangle and isolate the behavioral null-space, we rotate the $h$ vectors into an aligned space, using $Q_{\psi_{i}}$ and $Q_{\psi_{k}}$, where the behaviorally relevant information is organized along separate dimensions than the extraneous, behavioral null-space. We can then intervene on and transfer the relevant information without affecting other information. In the figure, this is done by applying binary masks (black represents 0s and white 1s and $\odot$ is a Hadamard product) to the $z$ vectors and taking their sum. We then invert the rotation on $z^{(v)}_{\psi_{i}}$ to return it to the original latent space where it can be used by $\psi_{i}$ to make predictions. (b) Depicts Stepwise MAS where the individual intervention shown in panel (a) is applied at multiple token positions in the sequence. We limit our interventions to contiguous sets of tokens starting from the first token and ending with a sampled position $t$.
@@
There are many existing methods for analyzing the similarity between different neural representations. Among these, there exist both correlational and causal methods. We see examples of the correlational variety in works that perform direct correlational analyses between individual ANN activations and BNN firing rates (Yamins & DiCarlo, 2016; Maheswaranathan et al., 2019; Khosla & Williams, 2023; Williams et al., 2022), and in works that use Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008)—or Centered Kernel Alignment (CKA) (Kornblith et al., 2019; Williams, 2024)—finding 2nd order isomorphisms between model and system. We also see examples of this in linear decoding techniques, where linear decodability and predictability can be used as a measure of the type of information encoded in distributed representations (Chen et al., 2020; Radford et al., 2021; Grill et al., 2020; Caron et al., 2021; Haxby et al., 2001; Haxby, 2013; Lampinen et al., 2024; Thobani et al., 2024). However, such correlative methods have received many criticisms for their sensitivity to noise structure (e.g. nuisance variables), reliance on outliers and large principal components, sensitivity to transformations that preserve functional behavior, and sensitivity to sample sizes and vector dimensionality (Dujmović et al., 2022; Popal et al., 2020; Davari et al., 2022; Murphy et al., 2024).
What do we miss when we ignore the functional, behavioral relevance of the neural activity? Many works have pointed out the disconnect between neural activity and its causal impact on network behavior (Pearl, 2010; Geiger et al., 2024; Cloos et al., 2024; Schaeffer et al., 2024; Lampinen et al., 2025). How do we incorporate a focus on behavioral relevance in our measures of neural similarity? Some works have emphasized network function in representational comparisons by transforming intermediate representations from one neural system into a usable form for another. One such method for performing such operations is known as model stitching. We see examples of this in ANN-BNN comparisons in works like Sexton & Love (2022) where they use transformed neural recordings in a trained vision model, and in ANN-ANN comparisons where stitching has been used both as a measure of functional representational similarity and for improving one of the two models (Lähner & Moeller, 2024; Moschella et al., 2023; Bansal et al., 2021; Lenc & Vedaldi, 2015; Klabunde et al., 2025; Braun et al., 2025). However, model stitching has its own limitations. Hernandez et al. (2022) showed that later model layers can successfully transfer behavior regardless of distance from the receiving layer, and Smith et al. (2025) showed that stitched representations can have lower intrinsic dimensionality than native representations, and networks trained on notably different tasks can be successfully stitched together.
@@
These shortcomings leave us with the broad question: how do we determine when neural systems are the same? In this work, we will pursue a notion of representational similarity that establishes an isomorphism between representational spaces with emphasis on preserving behavior. This means that we desire a mapping between model representations that is bidirectional (invertible), and we want the success of the mapping to be measured, at least in part, with respect to the networks’ function (i.e. behavior). In this work, we introduce such a method that uses causal interventions to isolate and compare behaviorally relevant activity from neural representations in different ANNs. The method has a close relationship to Distributed Alignment Search (DAS) (Geiger et al., 2021; 2023; Wu et al., 2024) for which we name it Model Alignment Search (MAS). MAS learns a rotation matrix, or more generally an alignment function, for each model in the comparison to simultaneously uncover causally relevant latent subspaces in each model and map these spaces onto one another so that behavioral information can be patched between models. The networks’ resulting behavior following an intervention can then be compared to the expected counterfactual behavior as a means of measuring functional similarity.
@@
    We introduce and motivate MAS by comparing it to correlative representational similarity methods demonstrating a need to supplement correlational with causal methods depending on the research goals.
@@
    Lastly, we introduce a counterfactual latent auxiliary loss objective that can be used to find behaviorally relevant alignments in cases where one of the two models is causally inaccessible (making the technique potentially relevant for comparisons between ANNs and BNNs).
@@ 2.1 Model Stitching
where $x\sim\mathcal{D}$ are samples from a dataset and $\ell$ is a divergence measure (e.g., KL divergence or cross-entropy) between $\psi_{1}$’s output distribution and $\psi_{2}$’s output distribution under the stitched representation.
@@
Model stitching has been used as a method to make claims about the similarity between two different NNs Bansal et al. (2021). However, it has problems as a similarity method due to the fact that it is unidirectional, and, as we will show in Section 4.2, it can make use of $\psi_{1}^{\prime}s$ behavioral null space when mapping to $\psi_{2}$—where the behavioral null space is defined as the span of vector directions that do not change the NN’s behavior. Furthermore, a successful mapping can ignore naturally occurring variability in $\psi_{2}$’s representations (Smith et al., 2025). Thus, a successful mapping $T$ only needs to produce a small set of sufficient representations for any given behavior meaning that successful $T$’s do not necessarily indicate that $\psi_{1}$ and $\psi_{2}$ perform the task in the same way.
@@ 2.2 DAS Formulation
DAS is a framework for causally determining the degree of alignment between a single NN’s latent representations and high-level variables from a causal abstraction (Geiger et al., 2021; 2023; Wu et al., 2024). We can think of such variables as those in a computer program, or nodes in a Directed Acyclic Graph (DAG), that can take on a well defined set of values. DAS does this alignment by attempting to transform model latent states $h\in\mathbb{R}^{d_{\psi}}$ into aligned vectors $z\in\mathbb{R}^{d_{\psi}}$ that consist of contiguous subspaces that have an analogous causal effect on behavior as the high-level, causal variables from a pre-specified causal abstraction. The transformation is performed by learnable, invertible Alignment Function (AF), $\mathcal{A}$, as follows: $z=\mathcal{A}(h)$ (Grant et al., 2025). We narrow our focus to orthogonal AFs: $\mathcal{A}(h)=Qh$ where $Q\in\mathbb{R}^{d_{\psi}\times d_{\psi}}$ is an orthogonal matrix. The purpose of this transformation is that it organizes causal subspaces into separated, contiguous dimensions in $z$. This allows for causal interventions on the values of specific subspaces—analogously to changing value of a variable in a computer program—without affecting the values of the other subspaces.
@@ 3.3 MAS Formulation
With these observations, we can generalize Equation 3 to the multi-model case for an intervention on the subspace for $\text{var}_{k}$:
$h_{\psi_{i}}^{(v)}=Q_{\psi_{i}}^{-1}((\mathcal{I}-D_{\psi_{i},\text{var}_{k}})Q_{\psi_{i}}h^{(trg)}_{\psi_{i}}+D_{\psi_{j},\text{var}_{k}}Q_{\psi_{j}}h^{(src)}_{\psi_{j}})$ (6)
Where $i$ and $j$ can be any model index in the set of all models considered, $Q_{\psi_{i}}$ is a scaled orthogonal rotation matrix for $\psi_{i}$, $D_{\psi_{i},\text{var}_{k}}\in R^{d_{\psi_{i}}\times d_{\psi_{i}}}$ is a diagonal, binary matrix, specific to each model, with $d_{\text{var}_{k}}$ non-zero elements used to isolate the dimensions corresponding to $\vec{z}_{\text{var}_{k}}$, $\mathcal{I}\in R^{d_{\psi_{i}}\times d_{\psi_{i}}}$ is the identity matrix, $h^{(src)}_{\psi_{j}}$ is the source vector from which the subspace is harvested, $h^{(trg)}_{\psi_{i}}$ is the target vector into which activity is substituted, and $h^{(v)}_{\psi_{i}}$ is the resulting intervened vector that replaces $h^{(trg)}_{\psi_{i}}$ in $\psi^{(trg)}_{i}$’s processing, allowing the model to make causally intervened predictions. In cases where $d_{\psi_{i}}>d_{\psi_{j}}$, $D_{\psi_{j}}$ can be expanded to shape $\mathbb{R}^{d_{\psi_{i}}\times d_{\psi_{j}}}$ using zero padding or reduced by removing rows from the bottom when $d_{\psi_{i}}<d_{\psi_{j}}$. See Figure 1(a) for a visualization.
For many of our analyses, we will perform MAS without segregating the behaviorally relevant subspace. Concretely, this can be formalized as a causal abstraction that uses a single Full variable in which all behaviorally relevant information is encoded, $\vec{z}_{full}$, and all extraneous information is encoded in $\vec{z}_{extra}$, the behavioral null-space and dormant subspaces. For completion: $z_{\psi_{i}}=\begin{bmatrix}\vec{z}_{\psi_{i},full}\\ \vec{z}_{\psi_{i},extra}\end{bmatrix}$, and we will freely isolate and manipulate $\vec{z}_{\psi_{i},full}$ using Equation 3 or  6.
@@ 3.4 MAS Training
It is crucial to include interventions where $i=j$ as this adds a soft constraint to the alignment that encourages the separation of $\vec{z}_{full}$ from $\vec{z}_{extra}$. This applies gradient pressure for the interventions to only use causally relevant information.
@@ 3.5 Mas Variants
Figure 2: (a) and (b) A comparison of MAS on the left axes and CKA and RSA on the right axes. We examine both the input embeddings and the hidden state vectors for models trained on the Multi-Object task. Dashed lines indicate the values for comparing individual models with themselves. (a) Results for GRUs compared to GRUs, where RSA can give low estimates of embedding similarity for different model seeds whereas MAS shows the high causal transfer we might expect. (b) Results for GRUs compared to 2-layer Transformers where we see a similar effect as (a) in the embeddings using RSA and we see a potential over-estimation of similarity of the hidden states in CKA and RSA. This over-estimation is with respect to causal transfer, as prior work has shown the transformers to use anti-Markovian solutions, where they recompute the relevant information at each step in the sequence. This is reflected in the low MAS IIA (Grant et al., 2025). (c) IIA comparing finetuned toxic and nontoxic LLMs using stepwise MAS. We can see that toxic models have higher IIA with themselves than with the nontoxic models. Notably, there is no significant difference for the nontoxic models compared to themselves. (d) Comparison of the IIA from DAS and MAS for different sizes of the aligned subspace, and model stitching with different rank transformation matrices.